OpenStudy (anonymous):

In the equation x^2 + mx + n = 0, m and n are integers. The only possible value for x is –3. What is the value of m ? The answer is 6, I just don't know why.

OpenStudy (anonymous):

x^2 + mx + n = (x + 3)^2 x^2 + mx + n = x^2 + 6x + 9

OpenStudy (anonymous):

How did you get the (x+3)^2?

OpenStudy (anonymous):

Sorry, the part with the question marks just a - in from of the 3

OpenStudy (anonymous):

*front